use inductive reasoning to describe each pattern
4,6,9,13......

Answer by tal3ntful(1)

What you do is first, figure out the problem
4,6,9,13...
6-4 =2
9-6 =3
13-9=4
It is accumilationg upward, so it is increasing by one.